# Enhancing Information Freshness: An AoI Optimized Markov Decision Process Dedicated In the Underwater Task

### v1

@@ Abstract
Ocean exploration utilizing autonomous underwater vehicles (AUVs) via reinforcement learning (RL) has emerged as a significant research focus. However, underwater tasks have mostly failed due to the observation delay caused by acoustic communication in the Internet of underwater things. In this study, we present an AoI optimized Markov decision process (AoI-MDP) to improve the performance of underwater tasks. Specifically, AoI-MDP models observation delay as signal delay through statistical signal processing, and includes this delay as a new component in the state space. Additionally, we introduce wait time in the action space, and integrate AoI with reward functions to achieve joint optimization of information freshness and decision-making for AUVs leveraging RL for training. Finally, we apply this approach to the multi-AUV data collection task scenario as an example. Simulation results highlight the feasibility of AoI-MDP, which effectively minimizes AoI while showcasing superior performance in the task. To accelerate relevant research in this field, the code for simulation will be released as open-source in the future.
@@ I Introduction
As a significant indicator evaluation the freshness of information, age of information (AoI) is proposed to measure the time elapsed at the receiver since the last information was generated until the most recent information is received [7]. And it has been verified to solve the severe delay caused by constantly sampling and transmitting observation information [8]. Central to this consensus is that minimizing AoI can enhance the freshness of information, thereby facilitating the efficiency of subsequent decision-making process in the presence of observation delay [8]. Currently, numerous studies have focused on optimizing AoI to aid decision-making in the context of land-based or underwater tasks. For example, Messaoudi et al. optimized vehicle trajectories relying on minimizing average AoI while reducing energy consumption [9]. Similarly, Lyu et al. leveraged AoI to assess transmission delay impacts on state estimation, improving performance under energy constraints [10]. These studies primarily aim to reduce AoI by improving the motion strategies of agents, without considering the impact of information update strategies on AoI. They assume that agents instantaneously perform the current action upon receiving the previous information. However, this zero-wait strategy has been shown to be suboptimal in scenarios with high variability in delay times [11]. Conversely, it has been demonstrated that introducing a waiting period before updating can achieve lower average AoI. This highlights the necessity of integrating optimized information update strategies into the underwater tasks.
@@ II-A AoI Optimized Markov Decision Process
Action Space $\boldsymbol{\mathcal{A}}$: the action space of the AoI-MDP consists of the tuple $a_{i}=(a_{i}^{{}^{\prime}},Z_{i})\in\boldsymbol{\mathcal{{A}^{{}^{\prime}}}}\times\boldsymbol{Z}$, where $a_{i}$ denotes the actions taken by the AUV, while $Z_{i}$ indicates the wait time between observing the environmental information and decision-making at time $i$. Through jointly optimizing the wait time $Z_{i}$ and action $a_{i}$, we aim to minimize the AoI, enabling the AUV’s decision-making policy to converge to an optimal level.
Reward Function $\boldsymbol{\mathcal{R}}$: the reward function $r_{i}^{{}^{\prime}}$ in standard MDP comprises elements with different roles, such as penalizing failures, promoting efficiency, and encouraging cooperation, etc. Here, we introduce the time-averaged AoI as a new component of the reward function. Thus, the updated reward function can be represented by the tuple $r_{i}=(r^{{}^{\prime}}_{i},-\bar{\Delta})$. And the time averaged AoI can be computed as follows:
$\bar{\Delta}=\frac{\sum_{i=2}^{\mathcal{N}}((2Y_{i-1}+Y_{i}+Z_{i})\times(Y_{i}+Z_{i}))}{2\times(\sum_{i=1}^{\mathcal{N}}Z_{i}+\sum_{i=1}^{\mathcal{N}}Y_{i})},$ (3)
where $\mathcal{N}$ is the length of information signal. Therefore, the time averaged AoI can be minimized through RL training.
@@ III-A Task Description and Settings
Since open-sour underwater tasks are scarce, we consider the scenario of a multi-AUV data collection task as a classic example to evaluate the feasibility and effectiveness of the AoI-MDP. This task utilizes RL algorithms to train AUVs to collect data of sensor nodes in the Internet of underwater things, encompassing multiple objectives, such as maximizing sum data rate and collision avoidance, while minimizing energy consumption, etc. For the remaining details and parameters of the task, please refer to the previous work [4].
@@ IV Conclusion
In this study, we propose AoI-MDP to improve the performance of underwater tasks. AoI-MDP models observation delay as signal delay through SSP, and includes this delay as a new component in the state space. Additionally, AoI-MDP introduces wait time in the action space, and integrate AoI with reward functions to achieve joint optimization of information freshness and decision making for AUVs leveraging RL for training. Simulation results highlight the feasibility, effectiveness and generalization of AoI-MDP over standard MDP, which effectively minimizes AoI while showcasing superior performance in the underwater task. The simulation code will be released as open-source to advance research in the future.

### v2

@@ Abstract
Ocean exploration utilizing autonomous underwater vehicles (AUVs) via reinforcement learning (RL) has emerged as a significant research focus. However, underwater tasks have mostly failed due to the observation delay caused by acoustic communication in the Internet of underwater things. In this study, we present an AoI optimized Markov decision process (AoI-MDP) to improve the performance of underwater tasks. Specifically, AoI-MDP models observation delay as signal delay through statistical signal processing, and includes this delay as a new component in the state space. Additionally, we introduce wait time in the action space, and integrate AoI with reward functions to achieve joint optimization of information freshness and decision-making for AUVs leveraging RL for training. Finally, we apply this approach to the multi-AUV data collection task scenario as an example. Simulation results highlight the feasibility of AoI-MDP, which effectively minimizes AoI while showcasing superior performance in the task. To accelerate relevant research in this field, we have made the simulation codes available as open-source1.
@@ I Introduction
As a significant indicator evaluation the freshness of information, age of information (AoI) is proposed to measure the time elapsed at the receiver since the last information was generated until the most recent information is received [7]. And it has been verified to solve the severe delay caused by constantly sampling and transmitting observation information [8]. Central to this consensus is that minimizing AoI can enhance the freshness of information, thereby facilitating efficiency of subsequent decision-making process in the presence of observation delay [8]. Currently, numerous studies have focused on optimizing AoI to aid decision-making in the context of land-based or underwater tasks. For example, Messaoudi et al. optimized vehicle trajectories relying on minimizing average AoI while reducing energy consumption [9]. Similarly, Lyu et al. leveraged AoI to assess transmission delay impacts on state estimation, improving performance under energy constraints [10]. These studies primarily aim to reduce AoI by improving motion strategies of agents, without considering the impact of information update strategies on AoI. They assume that agents instantaneously perform the current action upon receiving previous information. However, this zero-wait strategy has been shown to be suboptimal in scenarios with high variability in delay times [11]. Conversely, it has been demonstrated that introducing waiting time before updating can achieve lower average AoI. This highlights the necessity of integrating optimized information update strategies into underwater tasks.
@@ II-A AoI Optimized Markov Decision Process
Action Space $\boldsymbol{\mathcal{A}}$: the action space of the AoI-MDP consists of the tuple $a_{i}=(a_{i}^{{}^{\prime}},Z_{i})\in\boldsymbol{\mathcal{{A}^{{}^{\prime}}}}\times\boldsymbol{Z}$, where $a_{i}^{{}^{\prime}}$ denotes the actions taken by the AUV, while $Z_{i}$ indicates the wait time between observing the environmental information and decision-making at time $i$. Through jointly optimizing the wait time $Z_{i}$ and action $a_{i}$, we aim to minimize the AoI, enabling the AUV’s decision-making policy to converge to an optimal level.
Reward Function $\boldsymbol{\mathcal{R}}$: the reward function $r_{i}^{{}^{\prime}}$ in standard MDP comprises elements with different roles, such as penalizing failures, promoting efficiency, and encouraging cooperation, etc. Here, we introduce the time-averaged AoI as a new component of the reward function. Thus, the updated reward function can be represented by the tuple $r_{i}=(r^{{}^{\prime}}_{i},-\bar{\Delta})$. And the time-averaged AoI can be computed as follows:
$\bar{\Delta}=\frac{\sum_{i=1}^{\mathcal{N}}((2Y_{i-1}+Y_{i}+Z_{i-1})\times(Y_{i}+Z_{i-1}))+S_{0}}{2\times(\sum_{i=1}^{\mathcal{N}}Z_{i-1}+\sum_{i=1}^{\mathcal{N}}Y_{i}+Y_{0})},$ (3)
where $\mathcal{N}$ is the length of information signal, $S_{0}=0.5\times(2\Delta_{0}+Y_{0})\times Y_{0}$. Therefore, the time averaged AoI can be minimized through RL training.
@@ III-A Task Description and Settings
Since open-source underwater tasks are scarce, we consider the scenario of a multi-AUV data collection task as a classic example to evaluate the feasibility and effectiveness of the AoI-MDP. This task utilizes RL algorithms to train AUVs to collect data of sensor nodes in the Internet of underwater things, encompassing multiple objectives, such as maximizing sum data rate and collision avoidance, while minimizing energy consumption, etc. For the remaining details and parameters of the task, please refer to the previous work [4].
@@ IV Conclusion
In this study, we propose AoI-MDP to improve the performance of underwater tasks. AoI-MDP models observation delay as signal delay through SSP, and includes this delay as a new component in the state space. Additionally, AoI-MDP introduces wait time in the action space, and integrate AoI with reward functions to achieve joint optimization of information freshness and decision making for AUVs leveraging RL for training. Simulation results highlight the feasibility, effectiveness and generalization of AoI-MDP over standard MDP, which effectively minimizes AoI while showcasing superior performance in the underwater task. The simulation code has been released as open-source to facilitate future research.